# Advances in Appfl: A Comprehensive and Extensible Federated Learning Framework

### v1

@@ V-C Decentralized Federated Learning
Decentralized federated learning (DFL) is yet another variant of FL that eliminates the need for a central server to aggregate models. Instead, each node trains its local model using its own data, requests model parameters from neighboring clients, and aggregates these with its local model [81, 104]. Appfl supports DFL by implementing a DFL node agent that inherits most functionalities of both an FL client and server, enabling it to train local models and handle requests from neighboring clients. In this case study we set up DFL experiments with six nodes, where each node has three neighbors, as shown in Figure 12(a). Each node holds a heterogeneously partitioned MNIST dataset with six to eight classes and trains the model for 20 epochs. During each epoch, the node updates its model for 100 steps with a batch size of 64 using the Adam optimizer with a learning rate of 0.001, then aggregates its local model with those of its three neighbors. The experiment is repeated five times with different random seeds. Figure 12(b) presents the average validation accuracy and its standard deviation on the MNIST validation set across the training process for the six DFL nodes.
Figure 12: (a) Topology of the DFL experiments. (b) DFL validation accuracy for the DFL nodes, where the accuracy is evaluated after aggregating the local models of the neighbor DFL nodes.
VI Conclusion
In this paper we present the recent advancements in Appfl, a redesigned federated learning framework to simplify FL usage by offering comprehensive solutions to various challenges and to advance FL research through an easy-to-use, modular interface that facilitates the seamless integration of new algorithms. We demonstrate the capability and extensibility of Appfl by employing it to benchmark various FL aspects and provide case studies across different FL variants. Appfl is open-sourced under the MIT License, and we actively encourage contributions from the community.
In our future work we plan to incorporate more advanced privacy-enhancing technologies into the framework, such as secure multi-party computation, homomorphic encryption, and trusted execution environments, to further ensure the security of FL experiments. We will also focus on secure storage and deployment of the FL-trained models, enabling involved clients to efficiently and safely access the trained models for inference. Furthermore, Appfl will be used to backend a web platform that aims to provide federated learning as a service, streamlining the end-to-end process of AI model development for domain scientists, from data preparation through model training to model deployment.
@@ Acknowledgment
This research utilizes computing resources provided by the National Artificial Intelligence Research Resource (NAIRR) Pilot, supported by award NAIRR240008. We also gratefully acknowledge Amazon Web Services (AWS) for providing cloud computing credits that were used to assist with benchmarking efforts for this paper.

### v2

@@ V-C Decentralized Federated Learning
Decentralized federated learning (DFL) is another FL variant that eliminates the need for a central server. Instead, each node trains its local model, requests model parameters from neighboring clients, and aggregates these with its local model [46]. Appfl supports DFL by implementing a DFL node agent that inherits functionalities of both an FL client and server, enabling it to train local models and handle requests from neighboring clients. This case study sets up DFL experiments with six nodes, where each node has three neighbors, as shown in Figure 12(a). Each node holds a heterogeneously partitioned MNIST dataset with six to eight classes and trains the model for 20 epochs. During each epoch, the node updates its model for 100 steps with a batch size of 64 using the Adam optimizer with a learning rate of 0.001, then aggregates its local model with those of its three neighbors. The experiment is repeated five times. Figure 12(b) presents the average validation accuracy and its standard deviation on the MNIST validation set across the training process for the six DFL nodes.
Figure 12: (a) Topology of the DFL experiments. (b) DFL validation accuracy for the DFL nodes, evaluated after aggregating the local models of the neighbor DFL nodes.
VI Conclusion and Future Work
In this paper, we present the recent advancements in Appfl, a federated learning framework to simplify FL usage by offering comprehensive solutions to various challenges and to advance FL research through an easy-to-use, modular interface that facilitates the seamless integration of new algorithms. We demonstrate the capability and extensibility of Appfl by employing it to benchmark various FL components and provide case studies across different FL variants. Appfl is open-sourced under the MIT License, and we actively encourage contributions from the community. In our future work, we plan to incorporate more advanced privacy-enhancing technologies into the framework, such as secure multi-party computation, homomorphic encryption, and trusted execution environments, to further ensure the security of FL experiments. We also aim to employ Appfl for training larger-scale foundation models by leveraging private data from multiple data silos.
@@ Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357. This research utilizes computing resources provided by the National Artificial Intelligence Research Resource (NAIRR) Pilot, supported by award NAIRR240008. We also gratefully acknowledge Amazon Web Services (AWS) for providing cloud computing credits that were used to assist with benchmarking efforts for this paper.